The initial speed of a car having excellent brakes is 30 mi per hour. When the brakes are applied, the car stops at around 2.0 second. What is the car acceleration?


1
Expert's answer
2020-09-10T14:39:59-0400
"a=\frac{v}{t}\\a=\frac{30(0.44704)}{2}=6.7\frac{m}{s^2}"